### v1

# Enhancing Adversarial Robustness of Vision Language Models via Adversarial Mixture Prompt Tuning

Abstract
Large pre-trained Vision Language Models (VLMs) have excellent generalization capabilities but are highly susceptible to adversarial examples, presenting potential security risks. To improve the robustness of VLMs against adversarial examples, adversarial prompt tuning methods are proposed to align the text feature with the adversarial image feature without changing model parameters. However, when facing various adversarial attacks, a single learnable text prompt has insufficient generalization to align well with all adversarial image features, which finally leads to the overfitting phenomenon. To address the above challenge, in this paper, we empirically find that increasing the number of learned prompts can bring more robustness improvement than a longer prompt. Then we propose an adversarial tuning method named Adversarial Mixture Prompt Tuning (AMPT) to enhance the generalization towards various adversarial attacks for VLMs. AMPT aims to learn mixture text prompts to obtain more robust text features. To further enhance the adaptability, we propose a conditional weight router based on the input adversarial image to predict the mixture weights of multiple learned prompts, which helps obtain sample-specific aggregated text features aligning with different adversarial image features. A series of experiments show that our method can achieve better adversarial robustness than state-of-the-art methods on 11 datasets under different experimental settings.
1 Introduction
Large pre-trained Vision Language Models (VLMs) such as CLIP [27] have excellent generalization capabilities and can be regarded as foundation models [2] in different downstream tasks, e.g., image-text retrieval, zero-shot image classification, or image generation guidance. Due to its wide range of application scenarios, it places high requirements on security performance. However, despite its excellent performance, VLMs face many potential security risks [13, 24, 28], including the fact that visual models are vulnerable to adversarial examples [31], which can generate serious threat to the application in actual scenarios.
To eliminate this potential security risk, many works have been proposed to improve the robustness of VLMs to adversarial example, which can be mainly divided into two types, full-parameter fine-tuning [24, 32, 29, 35] and parameter-efficient fine-tuning [24, 32, 36, 17, 39, 21]. Among them, full-parameter fine-tuning is an effective method to improve the adversarial robustness of the model. However, this method often requires a lot of computational overhead and also affects the performance of the model on general tasks. Another type of parameter-efficient method, e.g., adversarial prompt tuning [36], freezes all or most of the weights of the model and only fine-tunes some of its parameters. This type of method can also effectively improve the adversarial robustness with lower training overhead compared with full-parameter fine-tuning, which is an ideal solution. However, adversarial prompt tuning may face a serious problem: insufficient generalization. For example, for the text prompt tuning [36, 17], when only one learnable prompt is fine-tuned, the text feature is not sufficient to fit the image features for various adversarial examples, which can easily lead to overfitting and further cause potential security risks [33].
Figure 1: The framework of Adversarial Mixture Prompt Tuning (AMPT). To enhance the adversarial robustness, we apply adversarial mixture prompt to generate diverse individual text feature, and utilize the conditional prompt weight router to obtain a sample-specific mixture text feature, and finally bring more generalization towards different adversarial examples.
To enhance the generalization of the adversarial text prompt, an intuitive approach is to increase the length of the text prompt. However, we find that when it grows to a certain length, a longer prompt will bring greater optimization difficulty, and also needs higher requirements on the corresponding text encoder to deal with long prompts, finally leading to suboptimal robustness. Inspired by the Mixture of Experts (MoE) paradigm [4], we adopt an alternative strategy: increasing the number of learnable base prompts. Similar to multiple experts in MoE, we construct a composite prompt by combining several base prompts, with their weights adaptively determined based on the characteristics of the adversarial image. This approach enables the generation of more diverse and expressive text features. Moreover, since each base prompt remains short and is easier to optimize, leading to improved adversarial robustness. A preliminary experiment has empirically verified our idea, that is, increasing prompt number can indeed bring stronger robustness than increasing prompt length.
Based on the above consideration, in this paper, we propose an adversarial prompt tuning method named Adversarial Mixture Prompt Tuning (AMPT) to enhance the adversarial robustness of VLMs. Specifically, we fix the parameters of the text and image encoders but only optimize adversarial mixture prompts. These text prompts pass through the text encoder and generate diverse individual text features. In addition, in order to enhance the adaptability, we propose a conditional text weight router based on image features to predict the weights of adversarial mixture prompts and aggregate them into a sample-specific aggregated text feature, so as to adaptively align with the diverse adversarial image features. A series of experiments show that our AMPT can achieve better accuracy and robustness than state-of-the-art methods on multiple different datasets. Meanwhile, AMPT also shows better generalization across different datasets. Our contribution can be summarized as follows:
  • We find that for adversarial text prompt tuning, increasing the number of learnable text prompts can achieve a better robustness improvement than only increasing the length of learnable text prompts within a certain range of parameters.
  • We propose a novel method named Adversarial Mixture Prompt Tuning (AMPT), which applies adversarial mixture prompts to generate diverse individual text features, where each text feature can play its unique roles for different adversarial examples, thereby alleviating the overfitting phenomenon.
  • We apply a conditional text weight router based on image features to predict the weights of different text features and obtain a sample-specific aggregated text feature that has pretty adaptability to align with different adversarial image features. Furthermore, we theoretically verify the effectiveness of the weight router.
  • We empirically verify the effectiveness of AMPT. Extensive experiments demonstrate that our AMPT can outperform state-of-the-art methods against adversarial examples in adversarial robustness and generalization across different datasets.
2 Related Work
2.1 Prompt Tuning for Accuracy in Vision-Language Models
Different from the methods of fine-tuning all model parameters, the prompt tuning method only fine-tunes the model’s input prompts. Through a training process, a learnable prompt suitable for downstream tasks is obtained to replace the hand-crafted prompt, thereby improving the performance of the VLMs. The prompt tuning methods are originated from text model [18, 19] and also have corresponding applications in visual models [14] and vision-language models [15, 38, 37]. CoOp [38] first utilizes a learnable vector to replace the hand-crafted in Vision-Language Models. Based on CoOp, CoCoOp [37] is proposed by introducing a conditional Meta-net based on an image feature to generate an instance-adaptive vector and add it to the learnable vector. Some research also tries to apply multiple prompts in VLMs [20, 6]. Different from the above works, in this paper, we mainly focus on improving the adversarial robustness via optimizing multiple prompts and embedding it into the adversarial training framework, which has obvious differences.
2.2 Adversarial Prompt Tuning in Vision-Language Models
Due to its excellent performance and low training cost, prompt tuning has been applied to improve the robustness of VLMs. [5] applies visual prompting to enhance the adversarial robustness. Furthermore, TeCoA [24] and PMG-AFT [32] employ visual prompt tuning to improve the adversarial robustness of VLMs. AdvPT [36] and APT [17] are proposed to apply text prompt tuning to further enhance the VLMs against image attacks. FAP [39] tries to enhance the robustness via bimodal tuning, while APD [21] further extends FAP into the adversarial distillation setting. To solve the insufficient generalization, [33] applies Test-Time Adversarial Prompt Tuning (TAPT) to learns defensive bimodal (textual and visual) prompts. Different from the above research, in this paper, we try to improve the adversarial robustness via adversarial mixture prompt tuning in the training process.
3 The Necessity of Adversarial Mixture Prompts
3.1 Formulation of Adversarial Prompt Tuning in CLIP
CoOp [38] first applies the text prompt tuning in CLIP to improve the performance of downstream tasks, and [36, 17] apply the adversarial prompt tuning in improving adversarial robustness, and the optimization goal of adversarial prompt tuning can be defined as follows:
$\mathop{\arg\min}_{t}\mathbb{E}_{(x,t,y)\sim\mathcal{D}}(\mathcal{L}(\tilde{x},t,y;F_{\theta_{v}},F_{\theta_{t}})),$ (1)
where $x$ and $t$ are the image and text pairs belong to the dataset $\mathcal{D}$. For the image classification task with $N$ classes, texts $t$ also contain $N$ different prompts: $\{t_{1},t_{2},\cdots,t_{N}\}$. $\tilde{x}$ denotes adversarial examples. $y$ denotes the ground truth. $y_{ij}$ indicates whether the image $x_{i}$ and text $t_{j}$ pair match, if the image $x_{i}$ and text $t_{j}$ match, $y_{ij}$ is equal to 1, otherwise $y_{ij}$ is equal to 0; the $F_{\theta_{v}}$ and $F_{\theta_{t}}$ are the image encoder and text encoder of CLIP.
Meanwhile, as for the text $t$, a fixed text template, e.g., "a photo of a [CLASS]", is often directly used as the text input, and the maximum similarity between it and the input image is calculated to determine which class the image belongs to. [36, 17] apply a learnable text prompt, which consists of the class context and a learnable context and can be defined as follows:
$t_{j}=[context_{front}][CLASS_{j}][context_{end}].$ (2)
The image feature $z_{v}^{i}$ is generated by image encoder $F_{\theta_{v}}$ of input $\tilde{x}_{i}$, the text feature $z_{t}^{i}$ is generated by text encoder $F_{\theta_{t}}$ of input $t_{j}$, which can be defined as follows:
$\tilde{z}_{v}^{i}=F_{\theta_{v}}(\tilde{x}_{i}),z_{t}^{j}=F_{\theta_{t}}(t_{j}).$ (3)
For the image classification task, Cross-Entropy loss is applied as the optimization function in APT [17], which can be defined as follows:
$\mathcal{L}(\tilde{x}_{i},t,,y_{i};F_{\theta_{v}},F_{\theta_{t}})=\sum_{j}^{N}-y_{ij}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{j}))}{\sum_{j}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{j})},$ (4)
where the cosine similarity is applied to measure the degree of alignment of the features, and applying the softmax operation can obtain the probability that the $\tilde{x}_{i}$ aligns with the $z_{t}$.
3.2 Selecting a Longer Prompt or More Prompts in APT?
APT [17] extends the CoOp framework to enhance the robustness of VLMs against adversarial attacks. However, a single text prompt has potential generalization problems: when faced with complex adversarial examples, its parameters may struggle to adapt to the change. Therefore, we attempt to explore how to enhance the generalization of adversarial text prompt tuning from the perspectives of increasing length and increasing number.
To compare those two approaches, we keep the total prompt parameters the same. For example, we use a learnable prompt of length 64 to compare the robustness of 4 learnable prompts of length 16. We conducted experiments on five datasets and the results can be viewed in Figure 2. And the experiments are based on APT [17].
From the experimental results, we are surprised to find that in the process of adversarial prompt tuning, increasing the prompt number is more effective than increasing the prompt length. Specifically, we find that using 2/4 learnable adversarial prompts of length 16 achieves better adversarial robustness compared to using a single prompt of length 32/64, with an average improvement of 3.88%/4.34%. Notably, this setup also leads to an average performance gain of 4.56%/6.43% on clean samples. Furthermore, increasing the prompt number further enhances adversarial robustness; When the prompt number increases from 2 to 4, adversarial robustness improves by an additional 0.34%. In contrast, merely increasing the prompt length yields no obvious improvement in adversarial robustness.
Figure 2: The performance of adversarial prompt tuning with different length and number on five datasets. “APT-Lm-Nk” denotes the APT with prompt length m and prompt number k. We find that increasing the prompt number can enhance more robustness than increasing the prompt length (i.e., solid lines show better performance than dotted lines).
We argue that after the total number of parameters reaches a certain level, continuing to increase the length of prompts will increase the difficulty of learning an ideal prompt, so it hardly enhances the adversarial robustness. On the contrary, shorter prompts are relatively easier to learn, and adversarial mixture prompts can generate more diverse text features, which have more possibility to align with different adversarial examples. Therefore, increasing the prompt number can further improve robustness compared with increasing prompt length.
4 Adversarial Mixture Prompt Tuning
4.1 Overall Framework
Based on the above findings, we argue that the generalization of adversarial robustness can be improved by adding adversarial mixture prompts. Therefore, we propose Adversarial Mixture Prompt Tuning (AMPT) to further improve the adversarial robustness of the VLMs. Here our optimization goal can be formulated as follows:
$\mathop{\arg\min}_{T_{m}}\mathbb{E}_{(x,t,y)\sim\mathcal{D}}(\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})),$ (5)
where $\mathcal{L}_{AMPT}$ denotes the optimization loss function of our AMPT, and $T_{m}$ denotes the adversarial mixture prompts, $F_{\theta_{w}}$ denotes the conditional prompt weight router. As for the adversarial examples $\tilde{x}$, we follow the “on-the-fly” setting in [17], where the attacker can access all the parameters of the VLMs including the adversarial mixture prompts but can only apply adversarial perturbations to the image $x$. And the adversarial examples $\tilde{x}$ can be formulated as follows:
$\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})),$ (6)
where $\epsilon$ denotes the maximum perturbation scale. It should be mentioned that the “on-the-fly” setting is closer to the adversarial examples in [31, 22]. For the evaluation against adversarial attacks, we also follow this type of setting as default.
4.2 Adversarial Mixture Prompts
Assume adversarial mixture prompts $T_{m}$ have $K$ total of prompts, which can be defined as follows:
$T_{m}=\{t^{1},t^{2},\cdots,t^{K}\},$ (7)
where $t^{k}$ denotes the $k$-th learnable adversarial text prompt, which includes $N$ class text prompt: $\{t_{1}^{k},t_{2}^{k},\cdots,t_{N}^{k}\}$. Following [38, 17], the $CLASS_{j}$ context in each $t^{k}$ is represented by a sequence of class-specific vectors, and the learnable contexts are defined in the word embedding space, then $t^{k}$ for class $j$ can be formulated as follows:
$t_{j}^{k}=[V]_{1,j}^{k}...[V]_{M,j}^{k}[C_{j}],$ (8)
where the $M$ denotes the max length of learnable context. The position of $[C_{j}]$ can also be adjusted. Following [17], we apply the end position as the default position. It should be mentioned that the learnable context can also apply two different modes: unified context (UC) and class-specific context (CSC), where unified context denotes that $t_{j}^{k}$ for different classes share the same learnable context, and class-specific context denotes that $t_{j}^{k}$ for different classes has their independent learnable context.
For adversarial mixture prompts, We first input different text prompts into the text encoder to obtain individual text features, then we aggregate these text features into a mixture text feature, which can be formulated as follows:
$z_{t}^{j,i}=\sum_{k}^{K}\tilde{w}_{k}^{i}F_{\theta_{t}}(t_{j}^{k}),$ (9)
where $\tilde{w}_{k}^{i}$ denotes the weights of adversarial prompt $t^{k}$ for the adversarial examples $\tilde{x}_{i}$, and $\tilde{w}_{k}^{i}$ is irrelevant to the class (not effected by $j$). $z_{t}^{j,i}$ denotes the aggregated text feature of $j$-th class for the adversarial examples $\tilde{x}_{i}$. In this way, we can obtain adversarial mixture prompts with pretty diversity through adversarial training to defend against different adversarial examples.
4.3 Conditional Prompt Weight Router
Algorithm 1 Training Process of AMPT
1:The train dataset $\mathcal{D}$, clean examples $x$ and adversarial examples $\tilde{x}$, ground truth $y$, text encoder $F_{\theta_{t}}$ and image encoder $F_{\theta_{v}}$, adversarial mixture prompts with random initialization $T_{m}=\{t^{1},t^{2},\cdots,t^{K}\}$, total class number $N$, condition prompt weight router $F_{\theta_{w}}$ with random initialization parameter $\theta_{w}$, the max training epochs $max$-$epoch$.
2:for 00 to $max$-$epoch$ do
3:     for $Every~{}minibatch(x,t,y)~{}in~{}\mathcal{D}$ do
4:          $\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Get adversarial examples $\tilde{x}$.
5:          $\{z_{t,1},\cdots,z_{t,K}\}=\{F_{\theta_{t}}(t^{1}),\cdots,F_{\theta_{t}}(t^{K})\}$.   //  Get each feature of adversarial prompt.
6:         for $each~{}x_{i}~{}in~{}x$ do
7:              $\tilde{z}_{v}^{i}=F_{\theta_{v}}(\tilde{x})$.   //  Get the image feature of adversarial examples $\tilde{x}_{i}$.
8:               $\tilde{w}^{i}=softmax(F_{\theta_{w}}(\tilde{z}_{v}^{i}))$.   //  Obtain the text prompt weights $\tilde{w}^{i}$ based on $\tilde{x}_{i}$.
9:               $\{z_{t}^{1,i},\cdots,z_{t}^{N,i}\}=\{\sum_{k}^{K}\tilde{w}_{k}^{i}z_{t,k}^{1},\cdots,\sum_{k}^{K}\tilde{w}_{k}^{i}z_{t,k}^{N}\}$.   //  Get mixture text features for $\tilde{x}_{i}$.
10:         end for
11:          $\theta_{w}=\theta_{w}-\eta\cdot\nabla_{\theta_{w}}\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Update prompt weight router.
12:          $T_{m}=T_{m}-\eta\cdot\nabla_{T_{m}}\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Update mixture prompts.
13:     end for
14:end for
Although adversarial mixture prompts can provide diverse adversarial text features, how to select those diverse features still needs to be solved when facing different adversarial examples. Therefore, we focus to adjust the weights: $w_{k}^{i}$ in Eq. (9). The simplest approach is to convert $w_{k}^{i}$ to $1/K$. To cover diverse image adversarial examples, we propose the conditional prompt weight router, which can generate the image-specific multiple weights for the different adversarial prompts.
Here we design a light-weight network containing two full connection layers as the conditional prompt weight router to predict prompt weights $\tilde{w}^{i}=\{\tilde{w}_{1}^{i},\cdots,\tilde{w}_{K}^{i}\}$ of image $x_{i}$. Initially, we obtain the adversarial image feature from the image encoder of VLMs $\tilde{z}_{v}^{i}$, then we apply the conditional prompt weight router to predict the different weights, which can be formulated as follows:
$\tilde{w}^{i}=softmax(F_{\theta_{w}}(\tilde{z}_{v}^{i})/\tau_{w}),$ (10)
where $F_{\theta_{w}}$ denotes the conditional prompt weight router, the $z_{v}^{i}$ denotes the feature generated by image encoder of image $x_{i}$. The softmax operation can keep the sum of the weights as 1. The $\tau_{w}$ is applied to control the adjustment strength of the generated weight, while a smaller $\tau_{w}$ denotes stronger adjustment strength, and a larger $\tau_{w}$ denotes weaker adjustment strength, when $\tau_{w}$ approaches infinity, it will degenerate into $1/K$. With the assistance of an adaptive weight router, we can finally more generalizable and representative mixture text feature based on the image features, to further improve the adaptability to defend against different adversarial examples. Meanwhile, we also provide the Theorem 1 about our conditional prompt weight router.
Theorem 1.
Assume there are multiple different adversarial text prompts $T_{m}=\{t^{1},t^{2},\cdots,t^{K}\}$, and the corresponding error risk of $k$-th text prompt $t^{k}$ for adversarial examples $\tilde{x}$ is $\mathcal{R}(\tilde{x},t^{k},y)$, and the normalized prompt weights $\tilde{w}=\{\tilde{w}_{1},\tilde{w}_{2},\cdots,\tilde{w}_{K}\}$ are optimized to minimize the error risk expectation of adversarial example $\tilde{x}$, we can obtain:
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}(\tilde{x},t^{k},y))<\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}(\tilde{x},t^{k},y)).$ (11)
The proof of Theorem 1 can be viewed in Appendix A.1. Theorem 1 shows that conditional prompt weights can bring the smaller error expectation of the adversarial examples compared with the average error expectation of the adversarial examples, which further demonstrates the necessity and effectiveness of our conditional prompt weight router.
Then the entire process of our AMPT can be viewed in Figure 1, and the optimization loss function $\mathcal{L}_{AMPT}$ can be defined as follows:
$\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y_{i};F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})=-\sum_{j}^{N}y_{ij}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{j,i}))}{\sum_{j}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{j,i}))},$ (12)
and the final training process can be viewed in Algorithm 1. It should be mentioned that to minimize the computational cost, we further decouple Eq. (9) and compute each text feature in advance for a minibatch. For each image, the final mixture text feature is obtained based on pre-computed text features without redundant calculation.
5 Experiments
5.1 Experimental Setting
Datasets. Following [38, 17], we mainly conduct our experiments on 11 high-resolution vision datasets: ImageNet [10], Caltech101 [11], OxfordPets [26], StanfordCars [16], Flowers102 [25], Food101 [3], FGVCAircraft [23], SUN397 [34], DTD [7], EuroSAT [12], and UCF101 [30].
Models. Following the setting in [17], we apply ViT-B/32 as our default selected backbone of image encoder, and select the model trained by a strong AT method TeCoA [24] as our default optimized weight. It should be mentioned that our method can easily transform into other CLIP-liked structures.
Baselines. Because our AMPT is a text prompt tuning method, we mainly compare our method with some similar state-of-the-art methods: Hand Engineered Prompts (HEP), AdvPT [36] and APT [17]. Similar to CoOp [38], we select two versions for AMPT and APT: unified context (UC) and class-specific context (CSC) following APT [17], while AdvPT only contain the UC setting. We change the setting of AdvPT into the setting of APT [17] for the sake of fair comparison. Since AMPT can be well combined with bimodal prompt tuning, e.g., FAP [39] and APD [21], we do not conduct additional comparisons about them.
Evaluation Metric. Here we mainly select two adversarial attacks, PGD attack [22] and AutoAttack [9]. If without additional claim, we set the maximum perturbation $\epsilon$ of adversarial attacks to 4/255. For the PGD attack, we apply 100 iterations with a step $\epsilon/4$ following [17]. Meanwhile, we employ an esemble attack, AutoAttack (AA) [9], which consists of four different attack methods: Auto-PGD (APGD), the Difference of Logits Ratio (DLR) attack, FAB-Attack [8], and the black-box Square Attack [1]. All the methods are evaluated on the entire test test if without additional instruction. For the evaluation of ImageNet against Autoattack, we select the 5000 test set to reduce the calculation overhead following [17], while conducting the AutoAttack on the entire test set is too expensive.
Table 1: Robustness performance(%) under all data training in the Unified Context (UC) setting.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 20.51 3.94 1.76 8.21 4.84 9.50 Clean 44.60 88.88 75.58 81.49 41.47 21.66 53.78 50.34 45.42 79.58 63.44 58.75 PGD 9.05 56.95 12.97 28.46 2.82 2.04 20.27 6.80 5.79 10.37 12.79 15.30 AdvPT-UC [36] AA 7.02 55.13 11.07 24.73 1.62 1.26 18.79 5.50 4.06 9.22 10.73 13.56 Clean 41.48 88.32 72.58 80.88 37.42 20.49 52.19 47.29 35.32 68.67 59.00 54.88 PGD 12.57 63.65 24.56 44.90 8.93 7.05 26.24 13.15 13.11 24.51 21.89 23.69 APT-UC [17] AA 8.16 61.01 16.43 38.61 3.92 3.33 22.40 8.06 7.32 29.79 16.39 19.58 Clean 42.30 87.38 72.72 82.34 45.17 20.58 53.43 51.48 38.98 68.19 60.27 56.62 PGD 12.62 65.03 24.78 46.81 12.16 7.56 28.49 13.92 13.34 37.70 21.99 25.85 MAPT-UC (ours) AA 8.18 62.39 16.57 41.21 6.01 3.21 25.41 9.91 8.03 34.56 17.29 21.16
Table 2: Robustness performance(%) under 16-shot training in the Unified Context (UC) setting.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 20.51 3.94 1.76 8.21 4.84 9.50 Clean 43.09 87.58 73.29 74.46 37.07 19.92 46.45 47.28 36.05 61.40 56.01 52.96 PGD 8.72 50.67 12.84 20.99 2.69 2.07 16.13 6.45 4.31 9.07 10.52 13.13 AdvPT-UC [36] AA 6.72 49.53 10.47 16.89 1.69 0.96 14.54 5.17 2.99 7.36 9.09 11.40 Clean 41.12 86.29 67.29 76.41 31.6 20.31 45.86 44.92 30.39 64.33 53.16 51.06 PGD 12.27 56.75 19.98 37.52 7.70 6.15 21.51 10.94 7.90 25.54 16.55 20.26 APT-UC [17] AA 7.88 53.43 13.46 32.20 3.37 2.64 18.91 6.88 4.17 16.68 12.74 15.67 Clean 41.20 87.14 70.62 76.53 38.54 19.29 47.75 47.32 30.73 57.06 54.09 51.84 PGD 12.38 57.69 20.69 37.79 9.29 6.63 22.99 11.08 8.46 24.72 18.61 20.94 MAPT-UC (ours) AA 7.84 55.05 13.27 32.28 5.07 2.52 20.04 7.75 4.60 19.35 14.14 16.54
[…]
5.2 Robustness Performance
We conduct a benchmark evaluation of our AMPT and baseline approaches. Table 1 and Table 2 present the performance of various text prompt methods in 11 datasets in both full-data and 16-shot training settings. Based on the results, AMPT not only demonstrates strong adversarial robustness under different attacks but also maintains competitive accuracy.
Specifically compared to AdvPT and APT, our approach consistently delivers superior robustness and stronger data efficiency. Under full-data training, AMPT-UC improves robustness over Baseline by an average of 2.16% and 1.58% (PGD/AA). Under the 16-shot training, AMPT-UC surpasses the Baseline by an average of 0.68% / 0.87% (PGD/AA), highlighting that the AMPT enhances APT’s performance under different types of attacks.
We also compare the two variants of APT-CSC and AMPT-CSC in Table 3. AMPT-CSC outperforms APT-CSC by an average of 0.27% / 0.23% (PGD/AA). This is likely because the CSC variant contains more parameters and therefore has a higher capacity for both APT-CSC and AMPT-CSC, and the insufficient generalization of APT-CSC will be somewhat alleviated, but AMPT-CSC still achieves better robustness, showing the effectiveness.
Table 3: Robustness performance(%) under all data training in Class-Specific Context (CSC) setting.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.74 89.49 72.74 87.82 49.04 31.44 55.56 51.00 42.62 69.77 64.47 59.43 PGD 11.85 65.44 24.56 55.30 12.92 9.00 27.54 13.74 14.71 38.32 24.16 27.05 APT-CSC[17] AA 8.26 63.41 17.77 50.14 6.75 5.01 24.35 9.57 9.04 33.40 19.69 22.49 Clean 38.85 88.56 72.58 86.40 45.28 27.84 55.50 51.83 42.82 69.79 62.86 58.39 PGD 11.83 66.82 25.73 53.92 12.39 9.30 29.20 14.01 14.71 38.77 23.82 27.32 MAPT-CSC(ours) AA 7.94 64.67 18.12 49.05 6.50 4.86 25.83 9.88 9.13 34.65 19.30 22.72
5.3 Generalization across Different Datasets
Table 4: Comparsion of Robust Generalization (%) between APT-UC and AMPT-UC cross 9 different datasets.
Metric APT-UC AMPT-UC(ours)
Clean 14.75 18.39
PGD 3.39 4.89
AA 2.66 4.14
Meanwhile, we test the robust generalization across different datasets. We apply the APT-UC and AMPT-UC trained on Caltech101 with all-data training setting as source models and evaluate the adversarial robustness in different datasets, including OxfordPets, OxfordFlowers, StanfordCars, FGVCAircraft, Sun397, DTD, Food101, EruoSAT, and UCF101, and the average results are reported in Table 4. From the result, we find that our AMPT-UC has better robust generalization compared with APT-UC. Specifically, Our AMPT-UC has a 1.5% and 1.78% robustness improvement against PGD and AA attacks compared with APT-UC, showing that our AMPT has better performance in dealing with diverse adversarial examples even in unseen classes. The detailed results can be viewed in Appendix A.2.
5.4 Ablation Study
To verify the effectiveness of AMPT, we conduct a set of ablation studies. We conduct the experiment in the Caltech101 dataset with the 16-shot training setting with class-specific context prompts.
Effects of Different Components. Here we try to conduct ablation experiments on various components of the experiment. We use the single adversarial cue fine-tuning method as the baseline and add adversarial mixture prompts without adding the conditional prompt weight router. Furthermore, we add our conditional prompt weight router. The experimental results are shown in Table 5.
Table 5: Ablation Study towards different components.
Component Clean PGD
Baseline 86.29 56.75
Baseline+Mixture 87.14 59.19
Baseline+Mixture+Router 87.01 60.12
The results can verify the effectiveness of each component in our AMPT. Only adding mixture prompts but without weight router can bring a 2.44% improvement compared with the baseline against PGD attack, while conditional prompt weight router can further bring a 0.93% improvement. This result shows that the feature diversity brought by adversarial mixture prompts and the adaptability brought by the conditional prompt weight router work together to improve the generalization of the VLMs to deal with adversarial examples.
Table 6: Ablation Study towards Selection of Prompt Number.
Prompt number Clean PGD
1 86.29 56.75
3 86.98 59.55
5 87.66 59.67
8 87.01 60.12
10 86.98 60.04
Selection of Prompt number. We explore the selection of prompt numbers. We select the following text prompt number of our AMPT as 1, 3, 5, 8, 10, 12, and the result can be viewed in Table 6. From the results, when the prompt number increases at the beginning (from number 1 to 3), the adversarial robustness of AMPT will obviously increase. However, when it further increases (from number 8 to 10), the robustness remains basically unchanged. It can be explained that as the prompt number increases, the difficulty of prompt optimization also increases. Thus, we select the prompt number 8 as the default setting. Considering the calculation overhead, we also select prompt number 3 for the partial dataset.
Table 7: Ablation Study towards Selection of Hyper-parameter $\tau$.
$\tau$ Clean PGD
0.3 86.94 59.59
0.5 87.26 59.71
0.7 87.01 60.12
0.8 86.97 60.04
1 87.14 59.67
Selection of Hyper-parameter $\tau$. The temperatures $\tau$ can control the adjustment strength of the conditional prompt weight router. While smaller $\tau$ means larger adjustment strength, larger $\tau$ means smaller adjustment strength. We select the following text prompt numbers of our AMPT as 0.3, 0.5, 0.7, 0.8, and 1, and the results can be found in Table 7. Based on the experimental results, we select the Hyper-parameter $\tau$ to 0.7 in our final setting.
6 Limitation
From the perspective of computational cost, as a parameter-efficient method, AMPT is lightweight and can be quickly optimized. However, our AMPT tuning introduces additional parameters and still has more computational cost during the training process. In addition, similar to the other parameter-efficient method, AMPT relies on pre-trained VLMs with pretty initial adversarial robustness, which needs to be further explored in the future.
7 Conclusion
In this work, we focused on the overfitting problem of adversarial prompt tuning, and found that simply increasing the length of the text prompt led to the learning difficulty while increasing the prompt number was more likely to improve the adversarial robustness of the VLMs. Based on the observation, we proposed an adversarial prompt tuning method named Adversarial Mixture Prompt Tuning (AMPT), which introduced adversarial mixture prompts to obtain more general text features to improve adversarial robustness. In addition, we proposed a conditional prompt weight router, which generated different prompt weights for different text prompts based on the characteristics of the input image to improve the adaptability of adversarial mixture prompts. Meanwhile, we theoretically proved the effectiveness of the conditional prompt weight router. A series of experimental results showed that our AMPT could achieve obvious improvements for the in-distribution robustness on different datasets and had well-transfer robustness across different datasets. In the future, AMPT can transform into different modals, e.g., visual prompts, and will be applied to more different downstream tasks.
Appendix A Appendix
A.1 The proof of Theorem 1
Here we assume each text prompt keeps unchanged and only optimize the prompt weights $\tilde{w}=\{\tilde{w}_{1},\tilde{w}_{2},\cdots,\tilde{w}_{K}\}$ from initial weights $\tilde{w}^{0}=\{1/K,1/K,\cdots,1/K\}$. Then as for any adversarial example $\tilde{x}^{{}^{\prime}}$, without loss generalization, we can assmue a corresponding text prompt $t^{k}$ exist error risk as follows:
$0<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{1},y)<\cdots<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k-1},y)<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k+1},y)<\cdots<\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{K},y).$ (13)
Thus based on the Gradient Descent operation, the expectation of corresponding weights can be formulated as follows:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{1}{K}-\eta\cdot\frac{\partial\frac{1}{K}\sum_{k}^{K}\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)}{\partial\tilde{w}_{k}},$ (14)
and here we assume that different weights are independent of each other, so we can obtain the $\mathbb{E}(\tilde{w}_{K}^{{}^{\prime}})$ as follows:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y).$ (15)
Then, we apply softmax to each $\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})$, and denote the result of applying softmax to $\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})$ is $\mathbb{E}(\tilde{w}_{k})$, then we get:
$\mathbb{E}(\tilde{w}_{k}^{{}^{\prime}})=\frac{e^{\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)}}{\sum_{j}^{K}e^{\frac{1}{K}-\eta\cdot\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{j},y)}}.$ (16)
For simplicity, we let $\mathcal{R}_{k}=\mathcal{R}(\tilde{x}^{{}^{\prime}},t^{k},y)$, so Eq. (16)It can be further simplified to:
$\mathbb{E}(\tilde{w}_{k})=\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}.$ (17)
Then the entire error expectation of adversarial examples can be formulated as follows:
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}_{k})=\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}_{k})+\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}.$ (18)
After substituting $\mathbb{E}(\tilde{w}_{k})$ into $\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}$, we have:
$\sum_{k}^{K}(\mathbb{E}(\tilde{w}_{k})-\frac{1}{K})*\mathcal{R}_{k}=\sum_{k}^{K}(\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}-\frac{1}{K})\cdot\mathcal{R}_{k}.$ (19)
We assume the following relation holds:
$\sum_{k}^{K}(\frac{e^{-\eta\cdot\mathcal{R}_{k}}}{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}-\frac{1}{K})\cdot\mathcal{R}_{k}<0,$ (20)
the above inequality is equivalent to the following inequality:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (21)
When $K=2$, without loss of generality, we assume that $\mathcal{R}_{1}<\mathcal{R}_{2}$, then we have:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}=\frac{e^{-\eta\cdot\mathcal{R}_{1}}-e^{-\eta\cdot\mathcal{R}_{2}}}{2}\cdot(\mathcal{R}_{1}-\mathcal{R}_{2}).$ (22)
Since $\mathcal{R}_{1}<\mathcal{R}_{2}$, So we have $\mathcal{R}_{1}-\mathcal{R}_{2}<0$ and $\frac{e^{-\eta\cdot\mathcal{R}_{1}}-e^{-\eta\cdot\mathcal{R}_{2}}}{2}>0$, which means when $K=2$, the assumption holds. Next, we assume that the assumption holds when $K=n$, and when $K=n+1$, without loss of generality, suppose $\mathcal{R}_{n+1}$ satisfies:
$\mathcal{R}_{n+1}<\mathcal{R}_{1}<\dots<\mathcal{R}_{n},$ (23)
then we have:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}$ $=\sum_{k}^{n+1}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n+1}\cdot\mathcal{R}_{k}.$ (24)
By the induction hypothesis, we have:
$\sum_{k}^{n+1}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}$ $=\sum_{k}^{n}e^{-\eta\cdot\mathcal{R}_{k}}\cdot\mathcal{R}_{k}+e^{-\eta\cdot\mathcal{R}_{n+1}}\cdot\mathcal{R}_{n+1}$ (25)
$<\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}\cdot\sum_{k}^{n}\mathcal{R}_{k}+e^{-\eta\cdot\mathcal{R}_{n+1}}\cdot\mathcal{R}_{n+1},$
meanwhile, we can have:
$\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n+1}\cdot\mathcal{R}_{k}$ $=\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k}+\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\mathcal{R}_{n+1}$ (26)
$+\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k}+\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\mathcal{R}_{n+1}.$
Substituting Eq. (26) and Eq. (25) into Eq. (24), we can obtain:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1})\cdot\mathcal{R}_{k}$ $<[(\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n(n+1)}\cdot\sum_{k}^{n}\mathcal{R}_{k}-\frac{e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\sum_{k}^{n}\mathcal{R}_{k})$ (27)
$+(\frac{n\cdot e^{-\eta\cdot\mathcal{R}_{n+1}}}{n+1}\cdot\mathcal{R}_{n+1}-\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n+1}\cdot\mathcal{R}_{n+1})]$
$=\frac{1}{n+1}\cdot(\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}-e^{-\eta\cdot\mathcal{R}_{n+1}})\cdot(\sum_{k}^{n}\mathcal{R}_{k}-n\cdot\mathcal{R}_{n+1}).$
Since we have previously assumed that $\mathcal{R}_{n+1}<\mathcal{R}_{1}<\dots<\mathcal{R}_{n}$, so we can deduce that:
$\sum_{k}^{n}\mathcal{R}_{k}-n\cdot\mathcal{R}_{n+1}>0,$ (28)
and:
$\frac{\sum_{j}^{n}e^{-\eta\cdot\mathcal{R}_{j}}}{n}-e^{-\eta\cdot\mathcal{R}_{n+1}}<0.$ (29)
Therefore, we can conclude that:
$\sum_{k}^{n+1}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{n+1}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (30)
Therefore, by mathematical induction, we can conclude that the following holds for any K:
$\sum_{k}^{K}(e^{-\eta\cdot\mathcal{R}_{k}}-\frac{\sum_{j}^{K}e^{-\eta\cdot\mathcal{R}_{j}}}{K})\cdot\mathcal{R}_{k}<0.$ (31)
Substituting Eq. (31) into Eq. (18), we obtain:
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}_{k})<\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}_{k}).$ (32)
Then the Theorem 1 is proved.
A.2 Details on Generalization cross Different Datasets
As shown in Table 8, we present the robust generalization performance across different target datasets when trained on Caltech101 as the source domain. The results show that our AMPT-UC consistently outperforms APT-UC in adversarial robustness across 8 datasets, and achieves higher accuracy on 6 datasets, indicating superior generalization ability and enhanced resilience against diverse attacks.
Table 8: Detailed results comparing the Robust Generalization (%) between APT-UC and AMPT-UC cross 9 different datasets based on Caltech101 adversarial prompts.
Method Metric OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 29.95 14.01 8.10 1.83 16.84 14.19 15.52 12.81 19.51 14.75 PGD 10.06 3.17 0.87 0.36 7.74 2.25 1.43 1.14 3.49 3.39 APT-UC AA 8.67 2.15 0.39 0.33 6.74 1.63 0.77 0.58 2.72 2.66 Clean 45.08 16.00 13.01 2.34 16.02 20.17 15.3 11.43 26.20 18.39 PGD 9.10 4.02 1.02 0.81 7.68 4.10 2.24 10.67 4.34 4.89 MAPT-UC AA 7.14 2.84 0.53 0.75 6.97 4.02 1.31 10.57 3.09 4.14
A.3 Trade-off between Accuracy and Robustness
As shown in Fig. 3, we compare the performance improvement per dataset of our adversarially-trained prompt over the standard-trained prompt for unified context. Most adversarially trained vision models tend to improve robustness at the expense of accuracy, and adversarially trained prompts also exhibit this trade-off, which is expected. More importantly, we observe that for most datasets, the gain in robustness outweighs the drop in accuracy. Specifically, AMPT improves adversarial robustness by an average of +7.60%, while incurring only a modest drop of -3.03% in accuracy. For instance, on OxfordPets, robustness increases significantly by +8.61%, with a slight gain of +0.23% in accuracy. These results suggest that our method achieves a relatively favorable trade-off between accuracy and robustness.
Figure 3: Trade-off between Accuracy and Robustness ($M=16$).